John Smith has a credit card that charges 15​% annual interest on the monthly average daily balance for the billing cycle. The current billing cycle has 30 days. For 14 days his balance was ​$2,536.42. For 7 days the balance was $1,536.42 9 days the balance was ​$1,835.63. Find the average daily balance. Find the amount of interest.(Round to the nearest cent as​ needed
Answer by Theo(13342)

balance of $2536.42 * 14 days equals a total of $35509.88.
balance of $1536.42 * 7 days equals a total of $10754.94.
balance of $1835.63 * 9 days equals a total of $16520.67.

total balance for 30 days is $35509.88 + $10754.94 + $16520.67 = $62785.49.

average daily balance for 30 days is $62785.49 / 30 = $2092.849667.

interest charged would be 15% / 100 / 365 * 30 * $2092.849667 = $25.80 rounded to the nearest penny.

the daily interest rate would be 15 / 100 / 365 = .0004109589041.

this rate is multiplied by $2092.849667 to get the daily finance charge of .8600752056.

the daily finance charge of .8600752056 is multiplied by 30 to get the finance charge of $25.80 rounded to two decimal places for the billing cycle.

this assumes no daily compounding.

if daily compounding is included, the formula is more complex and the interest charge for the billing cycle is slightly higher.

here's a reference.

https://www.nerdwallet.com/blog/credit-cards/how-credit-card-interest-calculated/

different banks use different methods so it's always good to check your bank's method before trying to do the calculations yourself.